Find an algebraic expression for sin[arccos(3x)]. PLEASE show the steps
1
Expert's answer
2012-10-30T11:39:27-0400
sin( t )=sqrt(1-cos^2 ( t ) )
sin(arccos(3x))=sqrt(1-cos^2 (arccos(3x)))=sqrt(1-(3x)^2)=sqrt(1-9x^2)